# Adaptive Multimodal Person Recognition: A Robust Framework for Handling Missing Modalities

### v2

@@ V-B Trimodal System
For bimodal evaluation, we similarly set the unused modality to zero, allowing the network to rely solely on the two remaining modalities. Since our model employs confidence-aware fusion, even in these constrained conditions, it can dynamically adjust its reliance on the available modalities based on their reliability. This structured evaluation allows us to quantify the robustness of each modality while validating the model’s ability to make accurate predictions even when some information is missing.
@@
Model Face-only Gesture-only Voice-only Face+Voice Face+Gesture Gesture+Voice Trimodal
@@
Condition Face Gesture Voice Face+Voice Face+Gesture Gesture+Voice Trimodal #Speakers #Utterances
@@
Setting Face Gesture Voice Face+Voice Face+Gesture Gesture+Voice Trimodal $\Delta_{\text{Top-1}}$
@@
Based on the results, the model achieved a Top-1 accuracy of 99.51% in the trimodal setup, showing 0.5% improvement over the 99.11% achieved by the face-voice bimodal scenario. This performance gain is attributed to the inclusion of the gesture modality.
Additionally, comparing the face-only configuration (96.42%) with the face-gesture configuration (97.36%) indicates that the gesture modality can enhance performance in visual scenarios. However, using gesture as an auxiliary input alongside voice resulted in a performance drop from 96.13% in the voice-only setup to 94.97% in the voice-gesture configuration. These findings suggest that the gesture modality is most beneficial when combined with the face modality, likely due to their shared visual nature and complementary information.
Table VI presents top-1 speaker identification accuracy across different modality combinations, separated by whether the training and testing samples come from the same or different recording sessions. The results highlight the impact of session variability on speaker recognition performance.
In the Single Session setting, all modalities perform exceptionally well, with face and gesture achieving over 99% accuracy and voice slightly lower at 97.52%. This indicates that under consistent conditions, unimodal cues are highly effective. However, in the multiple sessions condition, accuracy drops and most notably for gestures (down to 65.76%), suggesting that motion features are particularly sensitive to session changes.
V-C Ablation and Comparison
@@
The most substantial performance drops were observed when removing the Confidence module and Data Augmentation. The exclusion of the Confidence module resulted in a significant decrease of $6.8$ percentage points in trimodal accuracy, highlighting its critical role in dynamically weighting features based on their reliability. Furthermore, the removal of Data Augmentation led to an additional $3.4\%$ drop, with unimodal performance, particularly for gestures, falling to its lowest point ($58.6\%$). These findings suggest that while the fusion logic is vital, the model’s ability to generalize and handle noise is heavily dependent on robust training signals and the ability to evaluate the quality of incoming data.
@@ VI Conclusion
In this paper, we introduced a robust multimodal framework for person identification, integrating face, gesture, and voice modalities through an adaptive and confidence-driven fusion strategy. Our Trimodal and Bimodal systems leverage a unified architecture incorporating modality-specific processing, cross-attention mechanisms, gated fusion, and mistake correction modules. This design effectively addresses the challenge of modality loss, significantly outperforming conventional unimodal and late-fusion approaches.
@@
The insights derived from modality-specific performance analyses underline the complementary nature of face and voice modalities, while highlighting the session-dependent challenges associated with gesture features. Our confidence-aware fusion mechanism dynamically adapts to modality availability, ensuring consistent identification accuracy even in scenarios with significant modality degradation.

### v3

@@ V-C Trimodal System
For bimodal evaluation, we similarly set the unused modality to partially zeroed values and add noise, allowing the network to rely solely on the two remaining modalities. Since our model employs confidence-aware fusion, even in these constrained conditions, it can dynamically adjust its reliance on the available modalities based on their reliability. This structured evaluation allows us to quantify the robustness of each modality while validating the model’s ability to make accurate predictions even when some information is missing.
@@
Model Face-only motion-only Voice-only Face+Voice Face+motion motion+Voice Trimodal
@@
Condition Face motion Voice Face+Voice Face+motion motion+Voice Trimodal #Speakers #Utterances
@@
Setting Face motion Voice Face+Voice Face+motion motion+Voice Trimodal $\Delta_{\text{Top-1}}$
@@
Based on the results, the model achieved a Top-1 accuracy of 99.51% in the trimodal setup, showing 0.5% improvement over the 99.11% achieved by the face-voice bimodal scenario. This performance gain is attributed to the inclusion of the motion modality.
Additionally, comparing the face-only configuration (96.42%) with the face-motion configuration (97.36%) indicates that the motion modality can enhance performance in visual scenarios. However, using motion as an auxiliary input alongside voice resulted in a performance drop from 96.13% in the voice-only setup to 94.97% in the voice-motion configuration. These findings suggest that the motion modality is most beneficial when combined with the face modality, likely due to their shared visual nature and complementary information.
As shown in Table VI, the framework integrates the strengths of each modality, with face and motion reaching over 99% accuracy in single-session evaluations. These results demonstrate the efficacy of body motion cues. In multisession scenarios, body motion functions as a situational enhancer when fused with other modalities. By utilizing motion data as a supplement, the framework remains resilient even when session variability affects individual modalities.
V-D Ablation and Comparison
@@
The most substantial performance drops were observed when removing the Confidence module and Data Augmentation. The exclusion of the Confidence module resulted in a significant decrease of $6.8$ percentage points in trimodal accuracy, highlighting its critical role in dynamically weighting features based on their reliability. Furthermore, the removal of Data Augmentation led to an additional $3.4\%$ drop, with unimodal performance, particularly for motion, falling to its lowest point ($58.6\%$). These findings suggest that while the fusion logic is vital, the model’s ability to generalize and handle noise is heavily dependent on robust training signals and the ability to evaluate the quality of incoming data.
@@ VI Conclusion
In this paper, we introduced a robust multimodal framework for person identification, integrating face, motion, and voice modalities through an adaptive and confidence-driven fusion strategy. Our Trimodal and Bimodal systems leverage a unified architecture incorporating modality-specific processing, cross-attention mechanisms, gated fusion, and mistake correction modules. This design addresses the challenge of modality loss and outperforming conventional unimodal and late-fusion approaches.
@@
The insights derived from modality-specific performance analyses underline the complementary nature of face and voice modalities, while highlighting the session-dependent challenges associated with motion features. Our confidence-aware fusion mechanism dynamically adapts to modality availability, ensuring consistent identification accuracy even in scenarios with significant modality degradation.